### v1

# Hessian spectrum at the global minimum of the spherical pure-like mixed $p$-spin glasses

Abstract
We study the large $N$-dimensional limit of the Hessian spectrum at the global minimum of a subclass of spherical mixed $p$-spin models, extending the relevant results in the regime of topology trivialization.
1 Introduction
Let $S^{N-1}(\sqrt{N})$ be the sphere of radius $\sqrt{N}$ in $N$-dimensional Euclidean space. The Hamiltonian of the spherical mixed $p$-spin model is given by
$H_{N}(\bm{\sigma})=\sum_{p=2}^{\infty}\frac{\gamma_{p}}{N^{(p-1)/2}}\sum_{i_{1},\ldots,i_{p}=1}^{N}J_{i_{1},\ldots,i_{p}}^{(p)}\sigma_{i_{1}}\cdots\sigma_{i_{p}},\quad\bm{\sigma}=(\sigma_{1},\dots,\sigma_{N})\in S^{N-1}(\sqrt{N}),$ (1.1)
where $J_{i_{1},\ldots,i_{p}}^{(p)}$ are i.i.d. real standard Gaussian random variables and $\left\{\gamma_{p}\right\}_{p\geq 2}$ is a sequence of deterministic constants such that
$\sum_{p=2}^{\infty}2^{p}\gamma_{p}^{2}<\infty.$ (1.2)
Let us define
$\xi(x):=\sum_{p=2}^{\infty}\gamma_{p}^{2}x^{p}.$
Then direct calculation yields that the covariance of the Hamiltonian satisfies
$\mathbb{E}\left[H_{N}(\bm{\sigma})H_{N}\left(\bm{\sigma}^{\prime}\right)\right]=N\xi\left(R\left(\bm{\sigma},\bm{\sigma}^{\prime}\right)\right),$
where $R\left(\bm{\sigma},\bm{\sigma}^{\prime}\right):=\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}\sigma_{i}^{\prime}$ is the overlap between any two spin configurations $\bm{\sigma}$ and $\bm{\sigma}^{\prime}$. We fix the variance of $H_{N}$ by assuming $\xi(1)=\sum_{p=2}^{\infty}\gamma_{p}^{2}=1.$ If $\xi(x)=x^{p}$ for some $p\geq 3$, the model (1.1) is called the pure $p$-spin model. The assumption (1.2) ensures that the infinite sum in the model (1.1) converges almost surely and the Hamiltonian $H_{N}$ is almost surely smooth and Morse [AT07, Theorem 11.3.1], which guarantees the application of the Kac-Rice formula [AT07, theorem 12.1.1]. Denote the ground state energy in the spherical version of the mixed $p$-spin model by
$GS_{N}=\frac{1}{N}\min_{\bm{\sigma}\in S^{N-1}(\sqrt{N})}H_{N}(\bm{\sigma}),$
and the corresponding ground state by
$\bm{\sigma}^{*}=\operatornamewithlimits{\mathrm{argmin}}_{\bm{\sigma}\in S^{N-1}(\sqrt{N})}H_{N}(\bm{\sigma}).$
To study the ground state energy, the classical method is to study the free energy
$F_{N}(\beta)=\frac{1}{\beta N}\log\int_{S^{N-1}(\sqrt{N})}e^{-\beta H_{N}(\bm{\sigma})}\Lambda_{N}(\mathrm{d}\bm{\sigma}),$
where $\Lambda_{N}$ is the normalized surface measure on the sphere $S^{N-1}(\sqrt{N})$. It is well-known that the limiting free energy can be computed through the famous Parisi formula [talagrand2006free], [chen2013aizenman]. From here, it can be deduced (cf. [AC17, (2) and (3)]) that the sequence $GS_{N}$ also converges to some value $GS$ almost surely, which is usually called limiting ground state energy,
$GS:=\lim_{N\rightarrow\infty}GS_{N},\quad a.s.$
The Parisi formula for $GS$ through the variational representation of Parisi PDE was derived in [AC17, Theorem 1]. Meanwhile, based on the Crisanti-Sommers representation, the Parisi formula for the limiting ground state energy $GS$ was also established in [chen] and [2016dynamic].
To understand the landscape of the Hamiltonian $H_{N}(\bm{\sigma})$, Auffinger, Ben Arous, and Černý [ABC13] investigated the mean number of critical points of the spherical pure $p$-spin model. For any Borel set $B\in\mathbb{R}$, denote by $\operatorname{Crt}_{N,k}(B)$ the number of critical points of $H_{N}(\bm{\sigma})$ when the value of $H_{N}(\bm{\sigma})/N$ is restricted to $B$ with index $k$. In this pioneering work, they proved that
$\lim_{N\rightarrow\infty}\frac{1}{N}\log\left(\mathbb{E}\left\{\operatorname{Crt}_{N,k}((-\infty,u))\right\}\right)=\Theta_{p,k}(u),$
where the explicit expression of $\Theta_{p,k}(u)$ is given in [ABC13, (2.16)], which is commonly referred to as the complexity function of the pure $p$-spin model with index $k$. The function $\Theta_{p,k}(u)$ is known to be non-decreasing function. Let $-E_{0}(p)$ be the unique zero of function $\Theta_{p,0}(u)$, and define $E_{\infty}(p)=2\sqrt{\frac{p-1}{p}}$. Moreover, the second moment of $\operatorname{Crt}_{N}((-\infty,u))$ was calculated by [subag2017complexity], which showed that the critical points are concentrated around their means: for any $p\geq 3$ and $u\in\left(-E_{0}(p),-E_{\infty}(p)\right)$,
$\lim_{N\rightarrow\infty}\frac{\mathbb{E}\left\{\left(\operatorname{Crt}_{N}((-\infty,u))\right)^{2}\right\}}{\left(\mathbb{E}\left\{\operatorname{Crt}_{N}((-\infty,u))\right\}\right)^{2}}=1.$
As a corollary, it was deduced in [subag2017complexity, Appendix IV] that
$GS=-E_{0}(p),\quad a.s.$ (1.3)
This result was also obtained earlier in [ABC13, Theorem 2.12] for $p\geq 4$ even. A similar result [ABA13] was extended to the spherical mixed $p$-spin model, which states that for any open set $B\in\mathbb{R}$
$\lim_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{E}\mathrm{Crt}_{N,k}(B)=\sup_{u\in B}\theta_{k,\xi}(u),$
where $\theta_{k,\xi}(u)$ is explicitly given in [ABA13, (2.10)]. Setting $E_{\infty}^{\prime}(\xi)=\frac{2\xi^{\prime}(1)\sqrt{\xi^{\prime\prime}(1)}}{\xi^{\prime}(1)+\xi^{\prime\prime}(1)}$ as in [ABA13, (1.12)], the function $\theta_{k,\xi}(u)$ is strictly increasing on $\left(-\infty,-E_{\infty}^{\prime}(\xi)\right)$ and strictly decreasing on $\left(-E_{\infty}^{\prime}(\xi),\infty\right)$ with the unique maximum $\theta_{k,\xi}(-E_{\infty}^{\prime}(\xi))>0$. Denote by $-E_{0}(\xi)$ the smallest zero of $\theta_{0,\xi}(u)$. In Section 4 of their work, they claimed that $\xi(x)$ belonging to the $pure$-$like$ mixture (cf. Definition 1.1) class is equivalent to $E_{0}(\xi)=f_{1}$, where $f_{1}$ can be interpreted as the zero-temperature limit of one-step replica symmetry breaking ($1$-RSB) Parisi functional [ABA13, Propsition 6]. Also through a second moment calculation, Ben Arous, Subag and Zeitouni [BSZ20] proved that $GS=-E_{0}(\xi)$ for the $pure$-$like$ mixture models under Assumption I.
Recently, [Be22] considered the spherical mixed $p$-spin model on the unit sphere $S^{N-1}$ with external field
$H_{N}^{h}(\bm{\sigma})=H_{N}(\bm{\sigma})+Nh\mathbf{u}_{N}\cdot\bm{\sigma},$
where $h\geq 0$ is the strength of the external field and $\mathbf{u}_{N}\in S^{N-1}$ is a deterministic sequence. When $h^{2}>\xi^{\prime\prime}(1)-\xi^{\prime}(1)$, they showed that the Hamiltonian $H_{N}^{h}(\bm{\sigma})$ has trivial geometry, meaning that only critical points of $H_{N}^{h}$ are a single maximum and a single minimum. In this setting, they also found the convergence of the largest eigenvalues of the Hessian at global maximum [Be22, Theorem 1.1]. Indeed, the similar argument can be used to prove the convergence of the empirical spectrum measures of the Hessian at global maximum, with the help of a large deviation principle (LDP) for the spectrum measures of GOE matrices [benarous1997large]. The similar result was extended to locally isotropic Gaussian random fields on $\mathbb{R}^{N}$ [XZ22] provided the value of external field is large, which is previously known in the physics literature [FLD18].
Then a natural question arises: what happens in the case of a small external field. For the general spherical mixed $p$-spin models, this may be difficult to handle, since it is very challenging to find the unique minimizer of the Parisi functional for the limiting ground state energy [chen, Theorem 1], and therefore the value of $GS$. However, it is known that $GS=-E_{0}(p)$ [subag2017complexity, Appendix IV] (see also [chen, Propisition 3]) for the pure $p$-spin model without external fields ($h$=0) and $GS=-E_{0}(\xi)$ [BSZ20, Theorem 1.1] for the $pure$-$like$ mixture models also without external fields (under Assumption I). This means that the smallest zero of the complexity function exactly matches the corresponding $GS$. Together with the analysis of complexity function, we are able to study the behavior of the Hessian spectrum at global minimum $\bm{\sigma}^{*}$ in the thermodynamic limit $N\rightarrow\infty$ for these subclass models, which is the main focus of this paper. For more general models, whether this property still holds remains unknown and requires more detailed investigation in the future. Furthermore, the $full$ (cf. Definition 1.1 below) mixture models contradict this above property (see [ABA13, Corollary 4.1]), which may pose a significant obstacle to solving this problem. Let
$G\left(\xi^{\prime},\xi^{\prime\prime}\right):=\log\frac{\xi^{\prime\prime}(1)}{\xi^{\prime}(1)}-\frac{\left(\xi^{\prime\prime}(1)-\xi^{\prime}(1)\right)\left(\xi^{\prime\prime}(1)-\xi^{\prime}(1)+\xi^{\prime}(1)^{2}\right)}{\xi^{\prime\prime}(1)\xi^{\prime}(1)^{2}}.$
Definition 1.1.
A mixed model with covariance function $\xi(x)$ is called $pure$-$like$ if $G\left(\xi^{\prime},\xi^{\prime\prime}\right)>0$, $critical$ if $G\left(\xi^{\prime},\xi^{\prime\prime}\right)=0$, or $full$ if $G\left(\xi^{\prime},\xi^{\prime\prime}\right)<0$.
From the definition, one can verify that pure $p$-spin model belongs to the $pure$-$like$ mixture class for all $p\geq 3$. However, due to the trivial relation (2.4) in pure $p$-spin model, we treat the pure $p$-spin case separately from the $pure$-$like$ but not pure $p$-spin case. For the latter case, we require the following condition as stated in [BSZ20].
Assumption I.
Assume that $\xi(x)$ is mixed and $pure$-$like$, that $\frac{d^{2}}{dr^{2}}\Psi_{\xi}^{0}(0)<0$, and that the maximum of $\Psi_{\xi}^{0}(r)$ on the interval $[-1,1]$ is obtained uniquely at $r=0$.
The explicit expression of the function $\Psi_{\xi}^{0}(r)$ is defined in [BSZ20, (1.11)]. Since it is not the key ingredient of this paper, we do not provide it here.
Let $\mathrm{GOE}_{N}$ be an $N\times N$ matrix from the Gaussian Orthogonal Ensemble (GOE), i.e., $\mathrm{GOE}_{N}$ is a real symmetric matrix whose entries $(\mathrm{GOE}_{N})_{ij},i\leq j$ are independent centered Gaussian random variables with variance
$\mathbb{E}(\mathrm{GOE}_{N})_{ij}^{2}=\frac{1+\delta_{ij}}{N}.$
We denote by $\mathcal{P}(\mathbb{R})$ the space of probability measures on $\mathbb{R}$, equipped with the bounded Lipschitz metric for two probability measures $\mu,\nu\in\mathcal{P}(\mathbb{R})$
$d(\mu,\nu)=\sup\left\{\left\lvert\int f\mathrm{d}\mu-\int f\mathrm{d}\nu\right\rvert\colon\|f\|_{\infty}\leq 1,\|f\|_{L}\leq 1\right\},$ (1.4)
where $\|f\|_{\infty}$ and $\|f\|_{L}$ denote the $L^{\infty}$ norm and Lipschitz constant of the function $f$, respectively. The empirical spectrum measure of an $N\times N$ matrix $M$ is denoted by $L_{N}(M)=\frac{1}{N}\sum_{i=1}^{N}\delta_{\lambda_{i}(M)}$, where $\lambda_{i}(M)$ are the eigenvalues of $M$. We further use $\lambda_{\min}(M)$ to denote the smallest eigenvalue of matrix $M$. Let $\sigma_{c,r}$ denote the semicircle measure with center $c$ and radius $r$, which has the following density with respect to Lebesgue measure
$\sigma_{c,r}(\mathrm{d}x)=\frac{2}{\pi r^{2}}\sqrt{r^{2}-(x-c)^{2}}\,\mathrm{d}x.$
In particular, we denote the standard semicircle measure by $\sigma_{\rm sc}=\sigma_{0,2}$. To save space, we sometimes also use the shorthand notation $\xi^{\prime}=\xi^{\prime}(1)$ and $\xi^{\prime}=\xi^{\prime\prime}(1)$.
We organize the rest of the paper as follows. We prove the convergence of the empirical spectrum measures and smallest eigenvalues of the Hessian at global minimum for the spherical pure $p$-spin model in Section 2. Section 3 is dedicated to the analysis of spherical $pure$-$like$ mixture models.
2 Pure $p$-spin model
It is more convenient to work with this type of Gaussian process on the unit sphere $S^{N-1}$. For any $\bm{\sigma}\in S^{N-1}$, we define
$h_{N}(\bm{\sigma})=\frac{1}{\sqrt{N}}H_{N}(\sqrt{N}\bm{\sigma}).$ (2.1)
Then, $h_{N}(\bm{\sigma})$ is again a centered Gaussian process with
$\mathbb{E}\left[h_{N}(\bm{\sigma})h_{N}\left(\bm{\sigma}^{\prime}\right)\right]=\xi\left(\bm{\sigma}\cdot\bm{\sigma}^{\prime}\right).$
We endow the unit sphere $S^{N-1}$ with the standard Riemannian structure, induced by the Euclidean Riemannian metric on $\mathbb{R}^{N}$. We fix an orthonormal frame field $E=\left(E_{i}\right)_{i=1}^{N-1}$ on $S^{N-1}$, meaning that $\left(E_{i}(\bm{\sigma})\right)_{i=1}^{N-1}$ is a local orthonormal basis of $T_{\bm{\sigma}}S^{N-1}$ for any $\bm{\sigma}\in S^{N-1}$. We then define the spherical gradient and Hessian of $h_{N}(\bm{\sigma})$ as
$\nabla_{\mathrm{sp}}h_{N}(\bm{\sigma})=\left(E_{i}h_{N}(\bm{\sigma})\right)_{i=1}^{N-1},\quad\nabla^{2}_{\mathrm{sp}}h_{N}(\bm{\sigma})=\left(E_{i}E_{j}h_{N}(\bm{\sigma})\right)_{i,j=1}^{N-1}.$
Furthermore, let $\partial_{r}h_{N}(\bm{\sigma})$ denote the radial derivative of $h_{N}(\bm{\sigma})$ at $\bm{\sigma}\in S^{N-1}$, which can be understood as the directional derivative of $h_{N}(\bm{\sigma})$ in Euclidean space in the direction $\bm{\sigma}$. For the relationship between $h_{N}(\bm{\sigma})$ and its derivatives, we refer the readers to [Be22, Lemma 3.2].
For any open sets $B\subset\mathbb{R}$ and $D\subset\mathbb{R}$, we define the number of critical points when the value of Hamiltonian is restricted to $\sqrt{N}B$ and the value of radial derivative is restricted to $\sqrt{N}D$
$\operatorname{Crt}_{N}(B,D)=\#\left\{\bm{\sigma}\in S^{N-1}:\nabla_{\mathrm{sp}}h_{N}(\bm{\sigma})=0,\frac{1}{\sqrt{N}}h_{N}(\bm{\sigma})\in B,\frac{1}{\sqrt{N}}\partial_{r}h_{N}(\bm{\sigma})\in D\right\}.$ (2.2)
By the Kac-Rice formula [AT07, theorem 12.1.1], we have
$\begin{split}&\mathbb{E}\operatorname{Crt}_{N}(B,D)\\ &=\omega_{N}\varphi_{\nabla_{\mathrm{sp}}h_{N}(\bm{\sigma})}(0)\mathbb{E}\left\{\left\lvert\operatorname{det}(\nabla_{\mathrm{sp}}^{2}h_{N}(\bm{\sigma}))\right\rvert 1_{\left\{\frac{1}{\sqrt{N}}h_{N}(\bm{\sigma})\in B,\frac{1}{\sqrt{N}}\partial_{r}h_{N}(\bm{\sigma})\in D\right\}}\middle|\nabla_{\mathrm{sp}}h_{N}(\bm{\sigma})=0\right\},\end{split}$ (2.3)
where $\varphi_{\nabla_{\mathrm{sp}}h_{N}(\sigma)}(0)$ is the density of $\nabla_{\mathrm{sp}}h_{N}(\sigma)$ at 0, $\omega_{N}=2\pi^{N/2}/\Gamma(N/2)$ is the surface area of the $(N-1)$-dimensional unit sphere. For the pure $p$-spin model, since the covariance matrix of $h_{N}(\bm{\sigma})$ and $\partial_{r}h_{N}(\bm{\sigma})$ is degenerate, and thus we trivially have
$\partial_{r}h_{N}(\bm{\sigma})=ph_{N}(\bm{\sigma}),\quad a.s.$ (2.4)
Moreover, note that the three sets of random variables
$h_{N}(\bm{\sigma}),\quad\nabla_{\mathrm{sp}}h_{N}(\bm{\sigma}),\quad\nabla_{\mathrm{sp}}^{2}h_{N}(\bm{\sigma})+\partial_{r}h_{N}(\bm{\sigma})\mathbf{I}_{N-1}$
are mutually independent, where $\nabla_{\mathrm{sp}}h_{N}(\sigma)\sim N\left(0,p\mathbf{I}_{N-1}\right)$, and
$\sqrt{\frac{1}{(N-1)p(p-1)}}\left(\nabla_{\mathrm{sp}}^{2}h_{N}(\bm{\sigma})+\partial_{r}h_{N}(\bm{\sigma})\mathbf{I}_{N-1}\right)$ (2.5)
is an $(N-1)\times(N-1)$ normalized GOE matrix. By relations (2.4) and (2.5), conditioning on $\partial_{r}h_{N}(\bm{\sigma})=\sqrt{N}y$, we deduce from (2.3) that
$\begin{split}&\mathbb{E}\operatorname{Crt}_{N}(B,D)\\ &=C_{p,N}\int_{D}\exp\left\{-\frac{Ny^{2}}{2p^{2}}\right\}\mathbb{E}\left\{\left|\operatorname{det}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)p(p-1)}}\mathbf{I}_{N-1}\right)\right|1_{E}\left(\frac{y}{p}\right)\right\}\mathrm{d}y,\end{split}$ (2.6)
where the constant
$C_{p,N}=\frac{\sqrt{N}(N-1)^{(N-1)/2}(p-1)^{(N-1)/2}}{2^{(N-2)/2}p\Gamma(\frac{N}{2})}.$
The complexity function of (2.6) is obtained in [Be22, Lemma 6.2] (since the value of external field $h=0$, one can easily check that the supremum over $\gamma$ is at $\gamma=0$), which says that for any open sets $B$ and $D$
$\begin{split}\lim_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{E}\mathrm{Crt}_{N}(B,D)&=\sup_{y\in D,y/p\in B}\left[\frac{1}{2}\log(p-1)-\frac{(p-2)y^{2}}{4p^{2}(p-1)}+\Phi\left(\frac{y}{\sqrt{2p(p-1)}}\right)\right]\\ &=:\sup_{y\in D,y/p\in B}R(y),\end{split}$ (2.7)
[…]
$R(py)=\Theta_{p}(y),$ (2.8)
where $\Theta_{p}(y)$ is the complexity function of spherical pure $p$-spin model given in [ABA13, (2.15)]. Thus, $R(y)$ is non-decreasing and continuous on $\mathbb{R}$, just like $\Theta_{p}(y)$. On the other hand, the limiting ground state energy for pure $p$-spin model was computed by Chen and Sen in [chen, Proposition 3]. Together with equality (1.3), if $\xi(x)=x^{p}$ for $p\geq 3$, then
$GS=-E_{0}(p)=\frac{-\sqrt{p}}{\sqrt{z+1}}\left(1+\frac{z}{p}\right),$ (2.9)
where $z>0$ is the unique solution to
$\frac{1}{p}=\frac{1+z}{z^{2}}\log(1+z)-\frac{1}{z}.$ (2.10)
It follows that
$pE_{0}(p)>2\sqrt{p(p-1)}.$ (2.11)
Set $f(z)=\frac{\sqrt{p}}{\sqrt{z+1}}(p+z)$ for $z>0$. Then $f^{\prime}(z)=\frac{\sqrt{p}}{\sqrt{z+1}}-\frac{\sqrt{p}(p+z)}{(z+1)^{3/2}}$ and solving the equation $f^{\prime}(z)=0$ implies that $f(z)$ attains its unique minimizer at $z=p-2$ with $f(p-2)=2\sqrt{p(p-1)}$. However, $z=p-2$ is clearly not a solution to the equation (2.10), and thus (2.11) follows since $f(z)=pE_{0}(p)$ by (2.9).
Recall the bounded Lipschitz metric $d$ between two probability measures given in (1.4). We exhibit our result for the pure $p$-spin model.
Theorem 2.1 (Pure $p$-spin model).
Assume $\xi(x)=x^{p}$ for $p\geq 3$ fixed. Let $\bm{\sigma}^{*}$ denote the global minimum of $H_{N}(\bm{\sigma})$. Then we have convergence in probability
$\lim_{N\to\infty}d(L_{N-1}(\nabla^{2}_{\mathrm{sp}}H_{N}(\bm{\sigma}^{*})),\sigma_{c_{p},r_{p}})=0,$ (2.12)
$\lim_{N\rightarrow\infty}\lambda_{\min}\left(\nabla^{2}_{\mathrm{sp}}H_{N}\left(\bm{\sigma}^{*}\right)\right)=c_{p}-r_{p},$ (2.13)
where $c_{p}=pE_{0}(p)$ with $E_{0}(p)$ given by (2.9) and $r_{p}=2\sqrt{p(p-1)}$.
Remark 2.1.
By the different scaling on the domain (2.1), one has for any $\bm{\sigma}\in S^{N-1}$,
$\nabla^{2}_{\mathrm{sp}}H_{N}\left(\sqrt{N}\bm{\sigma}\right)=\nabla^{2}_{\mathrm{sp}}h_{N}\left(\bm{\sigma}\right).$
For any $\varepsilon>0$ and $\delta>0$, we define
$\mathcal{A}_{N,p}^{\varepsilon,\delta}=\#\left\{\bm{\sigma}\in S^{N-1}\colon\nabla_{\mathrm{sp}}h_{N}(\bm{\sigma})=0,\;\frac{1}{\sqrt{N}}h_{N}(\bm{\sigma})\in(-E_{0}(p)-\varepsilon,-E_{0}(p)+\varepsilon),\right.$
$L_{N-1}\left(\nabla^{2}_{\mathrm{sp}}h_{N}(\bm{\sigma})\right)\notin B\left(\sigma_{c_{p},r_{p}},\delta\right)\right\},$
$\mathcal{B}_{N,p}^{\varepsilon,\delta}=\#\left\{\bm{\sigma}\in S^{N-1}\colon\nabla_{\mathrm{sp}}h_{N}(\bm{\sigma})=0,\;\frac{1}{\sqrt{N}}h_{N}(\bm{\sigma})\in(-E_{0}(p)-\varepsilon,-E_{0}(p)+\varepsilon),\right.$
$\left|\lambda_{\min}\left(\nabla^{2}_{\mathrm{sp}}h_{N}(\bm{\sigma})\right)-c_{p}+r_{p}\right|>\delta\right\}.$
Lemma 2.2.
Assume $\xi(x)=x^{p}$ for some $p\geq 3$ fixed. Then for any $\delta>0$ fixed, there exists some $\varepsilon>0$ sufficiently small such that
$\lim_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{E}\mathcal{A}_{N,p}^{\varepsilon,\delta}=-\infty.$ (2.14)
Proof.
Using the Kac-Rice formula [AT07, theorem 12.1.1], similar to (2.6), we obtain
$\mathbb{E}\mathcal{A}_{N,p}^{\varepsilon,\delta}$
[…]
×$\quad\left.\times 1_{\left\{\frac{h_{N}(\bm{\sigma})}{\sqrt{N}}\in(-E_{0}(p)-\varepsilon,-E_{0}(p)+\varepsilon),\frac{\partial_{r}h_{N}(\bm{\sigma})}{\sqrt{N}}\in\mathbb{R},L_{N-1}\left(\nabla^{2}_{\mathrm{sp}}h_{N}(\bm{\sigma})\right)\notin B\left(\sigma_{c_{p},r_{p}},\delta\right)\right\}}\middle|\nabla_{\mathrm{sp}}h_{N}(\bm{\sigma})=0\right\}$
$=C_{p,N}\int_{-c_{p}-p\varepsilon}^{-c_{p}+p\varepsilon}\exp\left\{-\frac{Ny^{2}}{2p^{2}}\right\}\mathbb{E}\left\{\left|\operatorname{det}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)p(p-1)}}\mathbf{I}_{N-1}\right)\right|1_{A_{N}^{p}}\right\}\mathrm{d}y,$ (2.15)
where the event
$A_{N}^{p}:=\left\{L_{N-1}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)p(p-1)}}\mathbf{I}_{N-1}\right)\right.$
$\notin B\left(\sigma_{\frac{\sqrt{N}c_{p}}{\sqrt{(N-1)p(p-1)}},\frac{2\sqrt{N}}{\sqrt{N-1}}},\frac{\sqrt{N}\delta}{\sqrt{(N-1)p(p-1)}}\right)\right\}.$
Since $c_{p}=pE_{0}(p)>2\sqrt{p(p-1)}$ by (2.11), we may choose $\varepsilon>0$ to be sufficiently small such that for all $N$ large enough and any $y\in(-c_{p}-p\varepsilon,-c_{p}+p\varepsilon)$
$\frac{\sqrt{N}y}{\sqrt{(N-1)p(p-1)}}<-2.$ (2.16)
Note that for $|y+c_{p}|\leq p\varepsilon$ and a function $h$ with $\|h\|_{L}\leq 1$, for $N$ large enough
$\left\lvert\frac{1}{N-1}\sum_{i=1}^{N-1}\left[h\left(\lambda_{i}(\mathrm{GOE}_{N-1})-\frac{\sqrt{N}(y+c_{p})}{\sqrt{(N-1)p(p-1)}}\right)-h(\lambda_{i}\left(\mathrm{GOE}_{N-1})\right)\right]\right\rvert$
$\leq\frac{\sqrt{Np}\varepsilon}{\sqrt{(N-1)p-1}}\leq 2\varepsilon.$ (2.17)
Writing $c_{p}^{\prime}=\frac{c_{p}}{\sqrt{p(p-1)}}$, since $\frac{\sqrt{N}c_{p}}{\sqrt{(N-1)p(p-1)}}\rightarrow c_{p}^{\prime}$, $\frac{2\sqrt{N}}{\sqrt{N-1}}\rightarrow 2$ and $\frac{\sqrt{N}\delta}{\sqrt{(N-1)p(p-1)}}\rightarrow\frac{\delta}{\sqrt{p(p-1)}}$ as $N\rightarrow\infty$, it follows that
$d\left(\sigma_{\frac{\sqrt{N}c_{p}}{\sqrt{(N-1)p(p-1)}},\frac{2\sqrt{N}}{\sqrt{N-1}}},\sigma_{c_{p}^{\prime},2}\right)<\frac{\delta}{3\sqrt{p(p-1)}}.$
By the triangle inequality, for any $N$ large enough and $\varepsilon\leq\frac{\delta}{6\sqrt{p(p-1)}}$, if the event $A_{N}^{p}$ occurs, we deduce
$\mathrel{\phantom{=}}\frac{2\delta}{3\sqrt{p(p-1)}}$
$<d\left(L_{N-1}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)p(p-1)}}\mathbf{I}_{N-1}\right),\sigma_{c_{p}^{\prime},2}\right)$
$=d\left(L_{N-1}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}(y+c_{p})}{\sqrt{(N-1)p(p-1)}}\mathbf{I}_{N-1}\right),\sigma_{c_{p}^{\prime}-\frac{\sqrt{N}c_{p}}{\sqrt{(N-1)p(p-1)}},2}\right)$
$\leq d\left(L_{N-1}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}(y+c_{p})}{\sqrt{(N-1)p(p-1)}}\mathbf{I}_{N-1}\right),L_{N-1}(\mathrm{GOE}_{N-1})\right)$
$\quad+d(L_{N-1}(\mathrm{GOE}_{N-1}),\sigma_{\rm sc})+d\left(\sigma_{c_{p}^{\prime}-\frac{\sqrt{N}c_{p}}{\sqrt{(N-1)p(p-1)}},2},\sigma_{\rm sc}\right),$ (2.18)
which implies $d(L_{N-1}(\mathrm{GOE}_{N-1}),\sigma_{\rm sc})>\frac{\delta}{3\sqrt{p(p-1)}}=:\delta^{\prime}$ by (2). Together with the Cauchy–Schwarz inequality, we get for $N$ large enough
$\mathbb{E}\mathcal{A}_{N,p}^{\varepsilon,\delta}$
$\leq C_{p,N}\int_{-c_{p}-p\varepsilon}^{-c_{p}+p\varepsilon}\exp\left\{-\frac{Ny^{2}}{2p^{2}}\right\}$
×$\quad\times\mathbb{E}\left\{\left|\operatorname{det}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)p(p-1)}}\mathbf{I}_{N-1}\right)\right|1_{\left\{L_{N-1}\left(\mathrm{GOE}_{N-1}\right)\notin B\left(\sigma_{\rm{sc}},\delta^{\prime}\right)\right\}}\right\}\mathrm{d}y$
$\leq C_{p,N}\int_{-c_{p}-p\varepsilon}^{-c_{p}+p\varepsilon}\exp\left\{-\frac{Ny^{2}}{2p^{2}}\right\}\mathbb{P}\left(L_{N-1}\left(\mathrm{GOE}_{N-1}\right)\notin B\left(\sigma_{\rm{sc}},\delta^{\prime}\right)\right)^{\frac{1}{2}}$
×$\quad\quad\quad\times\left[\mathbb{E}\left\{\left|\operatorname{det}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)p(p-1)}}\mathbf{I}_{N-1}\right)\right|^{2}\right\}\right]^{\frac{1}{2}}\mathrm{d}y.$ (2.19)
By symmetry, property (2.16) allows us to use the upper bound for the second moment of the determinate of a shifted GOE [Be22, Lemma 6.4] (note the different scaling for the definition of standard GOE matrix). This implies that for large enough $N$ and small enough $\varepsilon$, it holds for any $\gamma>0$ and all $y\in(-c_{p}-p\varepsilon,-c_{p}+p\varepsilon)$ that
$\mathbb{E}\left\{\left|\operatorname{det}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)p(p-1)}}\mathbf{I}_{N-1}\right)\right|^{2}\right\}$
$\leq e^{\gamma N}\mathbb{E}\left\{\left|\operatorname{det}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)p(p-1)}}\mathbf{I}_{N-1}\right)\right|\right\}^{2}.$ (2.20)
Since the empirical spectral measures of GOE matrices satisfy an LDP with speed $N^{2}$ [benarous1997large], there exists a constant $C>0$ such that for $N$ large enough
$\mathbb{P}\left(L_{N-1}\left(\mathrm{GOE}_{N-1}\right)\notin B\left(\sigma_{\mathrm{sc}},\delta^{\prime}\right)\right)\leq e^{-CN^{2}},$ (2.21)
where $B(\sigma_{\mathrm{sc}},\delta^{\prime})$ denotes the open ball in the space $\mathcal{P}(\mathbb{R})$ with center $\sigma_{\mathrm{sc}}$ and radius $\delta^{\prime}$ with respect to the metric $d$ given by (1.4). Plugging the above two estimates (2) and (2.21) into (2) results in
$\mathbb{E}\mathcal{A}_{N,p}^{\varepsilon,\delta}$ $\leq C_{p,N}\int_{-c_{p}-p\varepsilon}^{-c_{p}+p\varepsilon}\exp\left\{-C^{\prime}N^{2}\right\}\exp\left\{-\frac{Ny^{2}}{2p^{2}}\right\}$
×$\quad\quad\quad\times\mathbb{E}\left\{\left|\operatorname{det}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)p(p-1)}}\mathbf{I}_{N-1}\right)\right|\right\}\mathrm{d}y,$
for some constant $C^{\prime}>0$ and all $N$ large enough. From (2.7), we obtain
$\mathbb{E}\mathcal{A}_{N,p}^{\varepsilon,\delta}\leq\exp\{-C^{\prime}N^{2}\}\exp\left\{\sup_{y\in(-c_{p}-p\varepsilon,-c_{p}+p\varepsilon)}NR(y)+o(N)\right\},$
which directly gives (2.14) by the continuity of $R(y)$ and the fact $R(-c_{p})=0$. ∎
Lemma 2.3.
Assume $\xi(x)=x^{p}$ for some $p\geq 3$ fixed. Then for any $\delta>0$ fixed, there exist some $\varepsilon>0$ sufficiently small and some $\eta>0$ such that
$\lim_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{E}\mathcal{B}_{N,p}^{\varepsilon,\delta}\leq-\eta.$ (2.22)
Proof.
The proof of this lemma follows the similar strategy as in Lemma 2.2. We need to address equality (2) with the event $A_{N}^{p}$ replaced by the event
$B_{N}^{p}:=\left\{\left|\lambda_{\min}\left(\mathrm{GOE}_{N-1}\right)-\frac{\sqrt{N}(y+c_{p})}{\sqrt{(N-1)p(p-1)}}+\frac{2\sqrt{N}}{\sqrt{N-1}}\right|>\frac{\sqrt{N}\delta}{\sqrt{(N-1)p(p-1)}}\right\}.$
For $y\in(-c_{p}-p\varepsilon,-c_{p}+p\varepsilon)$, once event $B_{N}^{p}$ occurs, for any $\delta>0$, we choose $\varepsilon>0$ to be sufficiently small and $N$ sufficiently large such that
$\left|\lambda_{\min}\left(\mathrm{GOE}_{N-1}\right)+2\right|>\frac{\delta}{2\sqrt{p(p-1)}},$
and thus we have
$\mathbb{E}\mathcal{B}_{N,p}^{\varepsilon,\delta}$
$\leq C_{p,N}\int_{-c_{p}-p\varepsilon}^{-c_{p}+p\varepsilon}\exp\left\{-\frac{Ny^{2}}{2p^{2}}\right\}$
×$\times\mathbb{E}\left\{\left|\operatorname{det}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)p(p-1)}}\mathbf{I}_{N-1}\right)\right|1_{\left\{\left|\lambda_{\min}\left(\mathrm{GOE}_{N-1}\right)+2\right|>\frac{\delta}{2\sqrt{p(p-1)}}\right\}}\right\}\mathrm{d}y.$ (2.23)
The LDP of the smallest eigenvalues of GOE matrices [BDG01] implies that for any $\delta>0$, there exists a $\zeta>0$ such that for $N$ large enough
$\mathbb{P}\left(\left|\lambda_{\min}\left(\mathrm{GOE}_{N}\right)+2\right|>\frac{\delta}{2\sqrt{p(p-1)}}\right)\leq e^{-\zeta N}.$ (2.24)
Note that the upper bound (2) holds for any $\gamma>0$. Here we take $\gamma<\frac{\zeta}{2}$. Hence, combining the Cauchy–Schwarz inequality, (2) and (2.24), we have for all $\varepsilon>0$ sufficiently small and $N$ large enough
$\mathbb{E}\mathcal{B}_{N,p}^{\varepsilon,\delta}$ $\leq C_{p,N}\int_{-c_{p}-p\varepsilon}^{-c_{p}+p\varepsilon}\exp\left\{-\frac{\zeta N}{2}\right\}\exp\left\{-\frac{Ny^{2}}{2p^{2}}\right\}$
×$\quad\quad\quad\times\mathbb{E}\left\{\left|\operatorname{det}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)p(p-1)}}\mathbf{I}_{N-1}\right)\right|\right\}\mathrm{d}y,$
which yields the desired result (2.22) for any $\eta<\frac{\zeta}{2}$ by the logarithmic asymptotics formula (2.7), the fact $R(-c_{p})=0$ and the continuity of $R(y)$. ∎
Now we are ready to prove Theorem 2.1.
Proof of Theorem 2.1.
Note that $GS=-E_{0}(p)$ from (2.9) for the pure $p$-spin model. To prove (2.12), it suffices to show that for any $\delta>0$,
$\mathbb{P}(L_{N-1}(\nabla^{2}_{\mathrm{sp}}H_{N}(\bm{\sigma}^{*}))\notin B(\sigma_{c_{p},r_{p}},\delta))\xrightarrow{N\rightarrow\infty}0.$
It is clearly that for any $\delta>0$ and $\varepsilon>0$
$\begin{split}&\mathrel{\phantom{=}}\mathbb{P}\left(L_{N-1}(\nabla^{2}_{\mathrm{sp}}H_{N}(\bm{\sigma}^{*}))\notin B(\sigma_{c_{p},r_{p}},\delta)\right)\\ &=\mathbb{P}\left(L_{N-1}(\nabla^{2}_{\mathrm{sp}}H_{N}(\bm{\sigma}^{*}))\notin B(\sigma_{c_{p},r_{p}},\delta),\frac{1}{N}H_{N}(\bm{\sigma}^{*})\in(-E_{0}(p)-\varepsilon,-E_{0}(p)+\varepsilon)\right)\\ &\quad+\mathbb{P}\left(L_{N-1}(\nabla^{2}_{\mathrm{sp}}H_{N}(\bm{\sigma}^{*}))\notin B(\sigma_{c_{p},r_{p}},\delta),\frac{1}{N}H_{N}(\bm{\sigma}^{*})\notin(-E_{0}(p)-\varepsilon,-E_{0}(p)+\varepsilon)\right).\end{split}$ (2.25)
Since $\frac{1}{N}H_{N}(\bm{\sigma}^{*})$ converges to $-E_{0}(p)$ almost surely as $N$ tends to $\infty$, the second term on the right hand side of (2.25) goes to 00 as $N$ tends to $\infty$ for any $\varepsilon>0$. If the event in the first term on the right hand side of (2.25) occurs, there exists at least one critical points $\bm{\sigma}\in S^{N-1}(\sqrt{N})$ with the restriction
$\frac{1}{N}H_{N}(\bm{\sigma})\in(-E_{0}(p)-\varepsilon,-E_{0}(p)+\varepsilon),\quad L_{N-1}(\nabla^{2}_{\mathrm{sp}}H_{N}(\bm{\sigma}))\notin B(\sigma_{c_{p},r_{p}},\delta).$
From the relation (2.1), the above event is equivalent to the event that there exists at least one critical points $\bm{\sigma}\in S^{N-1}$ with
$\frac{1}{\sqrt{N}}h_{N}(\bm{\sigma})\in(-E_{0}(p)-\varepsilon,-E_{0}(p)+\varepsilon),\quad L_{N-1}(\nabla^{2}_{\mathrm{sp}}h_{N}(\bm{\sigma}))\notin B(\sigma_{c_{p},r_{p}},\delta).$
Therefore, for any $\delta>0$ and $\varepsilon>0$, we have
$\mathbb{P}\left(L_{N-1}(\nabla^{2}_{\mathrm{sp}}H_{N}(\bm{\sigma}^{*}))\notin B(\sigma_{c_{p},r_{p}},\delta),\frac{1}{N}H_{N}(\bm{\sigma}^{*})\in(-E_{0}(p)-\varepsilon,-E_{0}(p)+\varepsilon)\right)$
$\leq\mathbb{P}(\mathcal{A}_{N,p}^{\varepsilon,\delta}\geq 1).$
However, Lemma 2.2 together with Markov’s inequality implies that for any $\delta>0$, there exists some $\varepsilon>0$ small enough such that
$\lim_{N\rightarrow\infty}\mathbb{P}(\mathcal{A}_{N,p}^{\varepsilon,\delta}\geq 1)=0,$
which completes the proof of (2.12). The above argument also can be applied to (2.13), since by Markov’s inequality, Lemma 2.3 directly gives $\lim_{N\rightarrow\infty}\mathbb{P}(\mathcal{B}_{N,p}^{\varepsilon,\delta}\geq 1)=0$. ∎
3 Pure-like Mixture models
Based on the covariance structure calculated in [Be22, Lemma 3.2], the three sets of random variables
$\left(h_{N}(\bm{\sigma}),\partial_{r}h_{N}(\bm{\sigma})\right),\quad\nabla_{\mathrm{sp}}h_{N}(\bm{\sigma}),\quad\nabla_{\mathrm{sp}}^{2}h_{N}(\bm{\sigma})+\partial_{r}h_{N}(\bm{\sigma})\mathbf{I}_{N-1}$
are mutually independent, where $\nabla_{\mathrm{sp}}h_{N}(\sigma)\sim N\left(0,\xi^{\prime}\left(1\right)\mathbf{I}_{N-1}\right)$, and
$\sqrt{\frac{1}{(N-1)\xi^{\prime\prime}\left(1\right)}}\left(\nabla_{\mathrm{sp}}^{2}h_{N}(\bm{\sigma})+\partial_{r}h_{N}(\bm{\sigma})\mathbf{I}_{N-1}\right)$ (3.1)
is an $(N-1)\times(N-1)$ GOE matrix and
$\left(h_{N}(\bm{\sigma}),\partial_{r}h_{N}(\bm{\sigma})\right)\sim N\left(0,\Sigma\right),$
with
$\Sigma:=\left(\begin{array}[]{cc}\xi\left(1\right)&\xi^{\prime}\left(1\right)\\ \xi^{\prime}\left(1\right)&\xi^{\prime\prime}\left(1\right)+\xi^{\prime}\left(1\right)\end{array}\right).$ (3.2)
From (3), conditioning on $\frac{h_{N}(\bm{\sigma})}{\sqrt{N}}=x$ and $\frac{\partial_{r}h_{N}(\bm{\sigma})}{\sqrt{N}}=y$, we immediately arrive at
$\mathbb{E}\operatorname{Crt}_{N}(B,D)$
$=C_{\xi,N}\int_{B\times D}\exp\left\{-\frac{N}{2}(x,y)\Sigma^{-1}(x,y)^{\top}\right\}$
×$\quad\quad\quad\times\mathbb{E}\left\{\left|\operatorname{det}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)\xi^{\prime\prime}\left(1\right)}}\mathbf{I}_{N-1}\right)\right|\right\}\mathrm{d}x\mathrm{d}y,$ (3.3)
where
$C_{\xi,N}=\frac{\sqrt{N}(N-1)^{(N-1)/2}\xi^{\prime\prime}(1)^{(N-1)/2}}{2^{(N-1)/2}\Gamma(\frac{N}{2})\xi^{\prime}(1)^{(N-1)/2}\sqrt{\pi\det\Sigma}}.$
Let us define
$\Psi_{*}(x)=\int_{\mathbb{R}}\log|x-t|\sigma_{\mathrm{sc}}(\mathrm{d}t).$
Direct calculation yields
$\Psi_{*}(x)=\begin{cases}\frac{x^{2}}{4}-\frac{1}{2}&\text{ if }0\leq|x|\leq 2,\\ \frac{x^{2}}{4}-\frac{1}{2}-\left[\frac{|x|}{4}\sqrt{x^{2}-4}-\log\left(\sqrt{\frac{x^{2}}{4}-1}+\frac{|x|}{2}\right)\right]&\text{ if }|x|>2.\end{cases}$ (3.4)
Based on $\Psi_{*}(x)$, we define
$F(x,y)$ $:=\frac{1}{2}+\frac{1}{2}\log\left(\frac{\xi^{\prime\prime}(1)}{\xi^{\prime}(1)}\right)-\frac{1}{2}(x,y)\Sigma^{-1}(x,y)^{\top}+\Psi_{*}\left(\frac{y}{\sqrt{\xi^{\prime\prime}(1)}}\right)$
$=\frac{1}{2}+\frac{1}{2}\log\left(\frac{\xi^{\prime\prime}(1)}{\xi^{\prime}(1)}\right)-\frac{x^{2}}{2}-\frac{\left(y-\xi^{\prime}(1)x\right)^{2}}{2\left(\xi^{\prime\prime}(1)+\xi^{\prime}(1)-\xi^{\prime}(1)^{2}\right)}+\Psi_{*}\left(\frac{y}{\sqrt{\xi^{\prime\prime}(1)}}\right).$ (3.5)
It was established in [BSZ20, Theorem 3.1] that for any intervals $B$ and $D$,
$\lim_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{E}\mathrm{Crt}_{N}(B,D)=\sup_{x\in B,y\in D}F(x,y).$ (3.6)
To maintain consistency in these series works, we define
$-E_{0}=\min\left\{x:\sup_{y\in\mathbb{R}}F(x,y)=0\right\},\quad-y_{0}=\operatornamewithlimits{argmax}_{y\in\mathbb{R}}F\left(-E_{0},y\right).$ (3.7)
Here we emphasize that for the $pure$-$like$ mixture models, $E_{0}$ is equal to $E_{0}(\xi)$ mentioned in the introduction, as claimed in footnote $2$ of [BSZ20]. The authors in [BSZ20, Theorem 1.1] proved that the sequence of ground state energy $GS_{N}$ converges to $-E_{0}$ almost surely under Assumption I, this is,
$\lim_{N\rightarrow\infty}GS_{N}=-E_{0},\quad a.s.$ (3.8)
We introduce the important threshold of energy level
$E_{\infty}=\frac{\xi^{\prime\prime}(1)+\xi^{\prime}(1)^{2}-\xi^{\prime}(1)}{\xi^{\prime}(1)\sqrt{\xi^{\prime\prime}(1)}}.$
From [BSZ20, Lemma 5.2], due to $-E_{0}<-E_{\infty}$, we know that if $\xi(x)$ belongs to the $pure$-$like$ mixture class, the solution $-y_{0}$ in (3.7) is unique and
$-y_{0}<-2\sqrt{\xi^{\prime\prime}(1)}.$ (3.9)
Later, we will prove that for the $pure$-$like$ mixture case, $y_{0}$ determines the center of the limiting empirical spectral measures of the Hessian at global minimum $\bm{\sigma}^{*}$.
Firstly, we present the main result of this section as below.
Theorem 3.1 (Pure-like mixture models).
Under Assumption I, let $\bm{\sigma}^{*}$ denote the global minimum of $H_{N}(\bm{\sigma})$. Then we have convergence in probability
$\lim_{N\to\infty}d(L_{N-1}(\nabla^{2}_{\mathrm{sp}}H_{N}(\bm{\sigma}^{*})),\sigma_{y_{0},2\sqrt{\xi^{\prime\prime}}})=0,$ (3.10)
$\lim_{N\rightarrow\infty}\lambda_{\min}\left(\nabla^{2}_{\mathrm{sp}}H_{N}\left(\bm{\sigma}^{*}\right)\right)=y_{0}-2\sqrt{\xi^{\prime\prime}},$ (3.11)
where $y_{0}$ is defined in (3.7).
Recall the function $F(x,y)$ defined in (3). By the definition of $-E_{0}$ and $-y_{0}$, since $-y_{0}<-2\sqrt{\xi^{\prime\prime}(1)}$ from (3.9), we know $F(-E_{0},-y_{0})=0$ and
$\frac{\partial F}{\partial y}(-E_{0},-y_{0})=-\frac{-y_{0}+\xi^{\prime}(1)E_{0}}{\xi^{\prime\prime}(1)+\xi^{\prime}(1)-\xi^{\prime}(1)^{2}}+\frac{-y_{0}+\sqrt{y_{0}^{2}-4\xi^{\prime\prime}(1)}}{2\xi^{\prime\prime}(1)}=0.$
We establish some properties of $F(x,y)$ when $x$ in a small neighborhood of $-E_{0}$, which is crucial for the subsequent proof.
Lemma 3.2.
If $\xi(x)$ is a pure-like mixture, then for sufficiently small $\varepsilon>0$, there exists some constant $\alpha>0$ such that
$F(x,y)<0,\quad\forall x\in(-E_{0}-\varepsilon,-E_{0}+\varepsilon),\;y\in\mathbb{R}\setminus(-y_{0}-\alpha\sqrt{\varepsilon},-y_{0}+\alpha\sqrt{\varepsilon}),$ (3.12)
and there exists some constant $\beta>0$ such that
$F(x,y)<\beta\sqrt{\varepsilon},\quad\forall x\in(-E_{0}-\varepsilon,-E_{0}+\varepsilon),\;y\in(-y_{0}-\alpha\sqrt{\varepsilon},-y_{0}+\alpha\sqrt{\varepsilon}).$ (3.13)
Proof.
[…]
$\Psi_{*}(x)$ $=\frac{x^{2}}{4}-\frac{1}{2}-\left[\frac{\lvert x\rvert}{4}\sqrt{x^{2}-4}-\log\left(\sqrt{\frac{x^{2}}{4}-1}+\frac{\lvert x\rvert}{2}\right)\right]$
$=-\frac{1}{2}+\frac{x^{2}-\lvert x\rvert\sqrt{x^{2}-4}}{4}+\log\left(\frac{\lvert x\rvert+\sqrt{x^{2}-4}}{2}\right)$
$=-\frac{1}{2}+\frac{\lvert x\rvert}{\lvert x\rvert+\sqrt{x^{2}-4}}-\log\left(\frac{\lvert x\rvert}{\lvert x\rvert+\sqrt{x^{2}-4}}\right)+\log\left(\frac{\lvert x\rvert}{2}\right).$
Since $t:=\lvert x\rvert/(\lvert x\rvert+\sqrt{x^{2}-4})\in(1/2,1)$, we have $t-\log t<1/2+\log 2$, which implies that
$\Psi_{*}(x)<-\frac{1}{2}+\frac{1}{2}+\log 2+\log\left(\frac{\lvert x\rvert}{2}\right)=\log\lvert x\rvert,\quad\text{for }\lvert x\rvert>2.$
If $\lvert x+E_{0}\rvert<\varepsilon$ and $y<\min\{-2\sqrt{\xi^{\prime\prime}(1)},-2E_{0}\xi^{\prime}(1)\}$, we have
$F(x,y)<\frac{1}{2}+\frac{1}{2}\log\left(\frac{\xi^{\prime\prime}(1)}{\xi^{\prime}(1)}\right)-\frac{(y+2\xi^{\prime}(1)E_{0})^{2}}{2(\xi^{\prime\prime}(1)+\xi^{\prime}(1)-\xi^{\prime}(1)^{2})}+\log\left\lvert\frac{y}{\sqrt{\xi^{\prime\prime}(1)}}\right\rvert.$
Therefore, for sufficiently small $\varepsilon>0$, there exists $K<-2\sqrt{\xi^{\prime\prime}(1)}$ sufficiently small, such that $F(x,y)<0$ for any $\lvert x+E_{0}\rvert<\varepsilon$ and $y<K$.
For $x\in(-E_{0}-\varepsilon,-E_{0}+\varepsilon)$, $y\in[K,0]$,
$F(x,y)-F(-E_{0},y)=-\frac{x^{2}-E_{0}^{2}}{2}-\frac{(y-\xi^{\prime}(1)x)^{2}-(y+\xi^{\prime}(1)E_{0})^{2}}{2(\xi^{\prime\prime}(1)+\xi^{\prime}(1)-\xi^{\prime}(1)^{2})}\leq C\varepsilon,$ (3.14)
where
$C=\frac{3E_{0}}{2}+\frac{-3K\xi^{\prime}(1)E_{0}+3\xi^{\prime}(1)^{2}E_{0}^{2}}{2(\xi^{\prime\prime}(1)+\xi^{\prime}(1)-\xi^{\prime}(1)^{2})}.$
It remains to estimate $F(-E_{0},y)$. Straightforward calculations show that
$\frac{\partial F}{\partial y}(-E_{0},y)=\begin{cases}-\frac{y+\xi^{\prime}(1)E_{0}}{\xi^{\prime\prime}(1)+\xi^{\prime}(1)-\xi^{\prime}(1)^{2}}+\frac{y}{2\xi^{\prime\prime}(1)},&-2\sqrt{\xi^{\prime\prime}(1)}\leq y\leq 0,\\ -\frac{y+\xi^{\prime}(1)E_{0}}{\xi^{\prime\prime}(1)+\xi^{\prime}(1)-\xi^{\prime}(1)^{2}}+\frac{y+\sqrt{y^{2}-4\xi^{\prime\prime}(1)}}{2\xi^{\prime\prime}(1)},&y<-2\sqrt{\xi^{\prime\prime}(1)}.\end{cases}$
which is a decreasing function on $(-\infty,0]$ with $\partial F/\partial y(-E_{0},-y_{0})=0$. For $y<-2\sqrt{\xi^{\prime\prime}(1)}$,
$\frac{\partial^{2}F}{\partial y^{2}}(-E_{0},y)=-c+\frac{y}{2\xi^{\prime\prime}(1)\sqrt{y^{2}-4\xi^{\prime\prime}(1)}}\leq-c,\quad$
where $c=\frac{\xi^{\prime\prime}(1)-\xi^{\prime}(1)+\xi^{\prime}(1)^{2}}{2\xi^{\prime\prime}(1)(\xi^{\prime\prime}(1)+\xi^{\prime}(1)-\xi^{\prime}(1)^{2})}>0$. Assume that $\varepsilon<(y_{0}-2\sqrt{\xi^{\prime\prime}(1)})^{2}$, then
$\frac{\partial F}{\partial y}(-E_{0},-y_{0}-\sqrt{\varepsilon})$ $=\frac{\partial F}{\partial y}(-E_{0},-y_{0}-\sqrt{\varepsilon})-\frac{\partial F}{\partial y}(-E_{0},-y_{0})\geq c\sqrt{\varepsilon},$
$\frac{\partial F}{\partial y}(-E_{0},-y_{0}+\sqrt{\varepsilon})$ $=\frac{\partial F}{\partial y}(-E_{0},-y_{0}+\sqrt{\varepsilon})-\frac{\partial F}{\partial y}(-E_{0},-y_{0})\leq-c\sqrt{\varepsilon}.$
Since $\partial F/\partial y(-E_{0},\cdot)$ is decreasing in $(-\infty,0]$, for $\alpha>1$, we have
$F(-E_{0},-y_{0}-\alpha\sqrt{\varepsilon})$ $=F(-E_{0},-y_{0}-\sqrt{\varepsilon})-\int_{-y_{0}-\alpha\sqrt{\varepsilon}}^{-y_{0}-\sqrt{\varepsilon}}\frac{\partial F}{\partial y}(-E_{0},y)\,\mathrm{d}y\leq-c(\alpha-1)\varepsilon,$
$F(-E_{0},-y_{0}+\alpha\sqrt{\varepsilon})$ $=F(-E_{0},-y_{0}+\sqrt{\varepsilon})+\int_{-y_{0}+\sqrt{\varepsilon}}^{-y_{0}+\alpha\sqrt{\varepsilon}}\frac{\partial F}{\partial y}(-E_{0},y)\,\mathrm{d}y\leq-c(\alpha-1)\varepsilon.$
Together with the monotonicity of $F(-E_{0},\cdot)$, we have
$F(-E_{0},y)\leq-c(\alpha-1)\varepsilon,\quad\forall y\in[K,0]\setminus(-y_{0}-\alpha\sqrt{\varepsilon},-y_{0}+\alpha\sqrt{\varepsilon}).$ (3.15)
[…]
For $y>0$, since $\Psi_{*}$ is an even function,
$F(x,y)=F(x,-y)+\frac{(y+\xi^{\prime}(1)x)^{2}-(y-\xi^{\prime}(1)x)^{2}}{2(\xi^{\prime\prime}(1)+\xi^{\prime}(1)-\xi^{\prime}(1)^{2})}.$
For $x<0$, it holds that $(y+\xi^{\prime}(1)x)^{2}<(y-\xi^{\prime}(1)x)^{2}$, which further implies that
$F(x,y)<F(x,-y)<0,\quad\forall x\in(-E_{0}-\varepsilon,-E_{0}+\varepsilon),\;y\in(0,+\infty)\setminus(y_{0}-\alpha\sqrt{\varepsilon},y_{0}+\alpha\sqrt{\varepsilon}).$
If $y\in(y_{0}-\alpha\sqrt{\varepsilon},y_{0}+\alpha\sqrt{\varepsilon})$,
$(y+\xi^{\prime}(1)x)^{2}-(y-\xi^{\prime}(1)x)^{2}=4\xi^{\prime}(1)xy\leq 4\xi^{\prime}(1)(-E_{0}+\varepsilon)(y_{0}-\alpha\sqrt{\varepsilon})<0.$
[…]
For $x\in(-E_{0}-\varepsilon,-E_{0}+\varepsilon)$, $y\in(-y_{0}-\alpha\sqrt{\varepsilon},-y_{0}+\alpha\sqrt{\varepsilon})$,
$F(x,y)-F(-E_{0},-y_{0})$ $=-\frac{x^{2}-E_{0}^{2}}{2}-\frac{(y-\xi^{\prime}(1)x)^{2}-(-y_{0}+\xi^{\prime}(1)E_{0})^{2}}{2(\xi^{\prime\prime}(1)+\xi^{\prime}(1)-\xi^{\prime}(1)^{2})}$
$\quad+\Psi_{*}\left(\frac{y}{\sqrt{\xi^{\prime\prime}(1)}}\right)-\Psi_{*}\left(\frac{-y_{0}}{\sqrt{\xi^{\prime\prime}(1)}}\right).$
Note that the function $\Psi_{*}$ is Lipschitz and denoted by $L_{\Psi_{*}}$ the Lipschitz constant. A little algebra yields that the claim (3.13) holds for sufficiently small $\varepsilon$ with
$\beta$ $>\frac{3\alpha(y_{0}+\xi^{\prime}E_{0})}{2(\xi^{\prime\prime}(1)+\xi^{\prime}(1)-\xi^{\prime}(1)^{2})}+\frac{\alpha}{\sqrt{\xi^{\prime\prime}}}L_{\Psi_{*}}.\qed$
Recall the definition of $-E_{0}$ and $-y_{0}$ in (3.7). For any $\varepsilon>0$ and $\delta>0$, we set
$\mathcal{A}_{N,\xi}^{\varepsilon,\delta}=\#\left\{\bm{\sigma}\in S^{N-1}\colon\nabla_{\mathrm{sp}}h_{N}(\bm{\sigma})=0,\;\frac{1}{\sqrt{N}}h_{N}(\bm{\sigma})\in(-E_{0}-\varepsilon,-E_{0}+\varepsilon),\right.$
$L_{N-1}\left(\nabla^{2}_{\mathrm{sp}}h_{N}(\bm{\sigma})\right)\notin B\left(\sigma_{y_{0},2\sqrt{\xi^{\prime\prime}}},\delta\right)\right\},$
$\mathcal{B}_{N,\xi}^{\varepsilon,\delta}=\#\left\{\bm{\sigma}\in S^{N-1}\colon\nabla_{\mathrm{sp}}h_{N}(\bm{\sigma})=0,\;\frac{1}{\sqrt{N}}h_{N}(\bm{\sigma})\in(-E_{0}-\varepsilon,-E_{0}+\varepsilon),\right.$
$\left|\lambda_{\min}\left(\nabla^{2}_{\mathrm{sp}}h_{N}(\bm{\sigma})\right)-y_{0}+2\sqrt{\xi^{\prime\prime}}\right|>\delta\right\}.$
Lemma 3.3.
If $\xi(x)$ is a pure-like mixture, then for any $\delta>0$ fixed, there exists some $\varepsilon>0$ sufficiently small such that
$\lim_{N\rightarrow\infty}\mathbb{E}\mathcal{A}_{N,\xi}^{\varepsilon,\delta}=0.$ (3.16)
Proof.
By the Kac-Rice formula [AT07, theorem 12.1.1], similar as to (3), we obtain
$\mathbb{E}\mathcal{A}_{N,\xi}^{\varepsilon,\delta}$
[…]
×$\quad\left.\times 1_{\left\{\frac{h_{N}(\bm{\sigma})}{\sqrt{N}}\in(-E_{0}-\varepsilon,-E_{0}+\varepsilon),\frac{\partial_{r}h_{N}(\bm{\sigma})}{\sqrt{N}}\in\mathbb{R},L_{N-1}\left(\nabla^{2}_{\mathrm{sp}}h_{N}(\bm{\sigma})\right)\notin B\left(\sigma_{y_{0},2\sqrt{\xi^{\prime\prime}}},\delta\right)\right\}}\middle|\nabla_{\mathrm{sp}}h_{N}(\bm{\sigma})=0\right\}$
$=C_{\xi,N}\int_{-E_{0}-\varepsilon}^{-E_{0}+\varepsilon}\int_{-\infty}^{\infty}\exp\left\{-\frac{N}{2}(x,y)\Sigma^{-1}(x,y)^{\top}\right\}$
×$\quad\quad\quad\times\mathbb{E}\left\{\left|\operatorname{det}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)\xi^{\prime\prime}}}\mathbf{I}_{N-1}\right)\right|1_{A_{N}^{\xi}}\right\}\mathrm{d}x\mathrm{d}y,$ (3.17)
where the event
$A_{N}^{\xi}=\left\{L_{N-1}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)\xi^{\prime\prime}}}\mathbf{I}_{N-1}\right)\notin B\left(\sigma_{\frac{\sqrt{N}y_{0}}{\sqrt{(N-1)\xi^{\prime\prime}}},\frac{2\sqrt{N}}{\sqrt{N-1}}},\frac{\sqrt{N}\delta}{\sqrt{(N-1)\xi^{\prime\prime}}}\right)\right\}.$
Based on the result of Lemma 3.2, we divide the integral (3) into three parts with respect to variable $y$
$\mathbb{E}\mathcal{A}_{N,\xi}^{\varepsilon,\delta}$ $=C_{\xi,N}\int_{-E_{0}-\varepsilon}^{-E_{0}+\varepsilon}\int_{-y_{0}-\alpha\sqrt{\varepsilon}}^{-y_{0}+\alpha\sqrt{\varepsilon}}\exp\left\{-\frac{N}{2}(x,y)\Sigma^{-1}(x,y)^{\top}\right\}$
×$\quad\quad\quad\times\mathbb{E}\left\{\left|\operatorname{det}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)\xi^{\prime\prime}}}\mathbf{I}_{N-1}\right)\right|1_{A_{N}^{\xi}}\right\}\mathrm{d}x\mathrm{d}y$
$\quad+C_{\xi,N}\int_{-E_{0}-\varepsilon}^{-E_{0}+\varepsilon}\int_{-\infty}^{-y_{0}-\alpha\sqrt{\varepsilon}}\exp\left\{-\frac{N}{2}(x,y)\Sigma^{-1}(x,y)^{\top}\right\}$
×$\quad\quad\quad\times\mathbb{E}\left\{\left|\operatorname{det}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)\xi^{\prime\prime}}}\mathbf{I}_{N-1}\right)\right|1_{A_{N}^{\xi}}\right\}\mathrm{d}x\mathrm{d}y$
$\quad+C_{\xi,N}\int_{-E_{0}-\varepsilon}^{-E_{0}+\varepsilon}\int_{-y_{0}+\alpha\sqrt{\varepsilon}}^{\infty}\exp\left\{-\frac{N}{2}(x,y)\Sigma^{-1}(x,y)^{\top}\right\}$
×$\quad\quad\quad\times\mathbb{E}\left\{\left|\operatorname{det}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)\xi^{\prime\prime}}}\mathbf{I}_{N-1}\right)\right|1_{A_{N}^{\xi}}\right\}\mathrm{d}x\mathrm{d}y$
$=:J_{N}^{1}+J_{N}^{2}+J_{N}^{3}.$ (3.18)
By the logarithmic asymptotics estimate (3.6) for the mean number of critical points, removing the indicator functions in $J_{N}^{2}$ and $J_{N}^{3}$ results in
$\limsup_{N\rightarrow\infty}\frac{1}{N}\log\left(J_{N}^{2}+J_{N}^{3}\right)\leq\sup_{\begin{subarray}{c}x\in(-E_{0}-\varepsilon,-E_{0}+\varepsilon)\\ y\in\mathbb{R}\setminus(-y_{0}-\alpha\sqrt{\varepsilon},-y_{0}+\alpha\sqrt{\varepsilon})\end{subarray}}F(x,y).$
By the property (3.12) in Lemma 3.2, there exists a sufficiently small $\varepsilon$ such that the supremum of the right hand side of above inequality is strictly smaller than 00. We then directly deduce for sufficiently small $\varepsilon$
$\lim_{N\rightarrow\infty}(J_{N}^{2}+J_{N}^{3})=0.$ (3.19)
It remains to consider $J_{N}^{1}$. Note that $-y_{0}<-2\sqrt{\xi^{\prime\prime}(1)}$ by (3.9). The fact that $\alpha\sqrt{\varepsilon}$ tends to 00 as $\varepsilon$ goes to 00 allows us to use the similar arguments as those in (2) and (2). We omit some tedious but straightforward steps to get for any $\varepsilon$ small enough and $N$ large enough
$J_{N}^{1}$ $\leq C_{\xi,N}\int_{-E_{0}-\varepsilon}^{-E_{0}+\varepsilon}\int_{-y_{0}-\alpha\sqrt{\varepsilon}}^{-y_{0}+\alpha\sqrt{\varepsilon}}\exp\left\{-\frac{N}{2}(x,y)\Sigma^{-1}(x,y)^{\top}\right\}$
×$\quad\times\mathbb{E}\left\{\left|\operatorname{det}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)\xi^{\prime\prime}}}\mathbf{I}_{N-1}\right)\right|1_{\left\{L_{N-1}\left(\mathrm{GOE}_{N-1}\right)\notin B\left(\sigma_{\rm{sc}},\frac{\delta}{3\sqrt{\xi^{\prime\prime}}}\right)\right\}}\right\}\mathrm{d}x\mathrm{d}y.$
Using the Cauchy–Schwarz inequality and replacing $p(p-1)$ by $\xi^{\prime\prime}$ in estimates (2) and (2.21), there exists some constant $C>0$ such that for any $\varepsilon>0$ sufficiently small and $N$ large enough
$J_{N}^{1}$ $\leq C_{\xi,N}\int_{-E_{0}-\varepsilon}^{-E_{0}+\varepsilon}\int_{-y_{0}-\alpha\sqrt{\varepsilon}}^{-y_{0}+\alpha\sqrt{\varepsilon}}\exp\left\{-CN^{2}\right\}\exp\left\{-\frac{N}{2}(x,y)\Sigma^{-1}(x,y)^{\top}\right\}$
×$\quad\times\mathbb{E}\left\{\left|\operatorname{det}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)\xi^{\prime\prime}}}\mathbf{I}_{N-1}\right)\right|\right\}\mathrm{d}x\mathrm{d}y.$
From here, we apply (3.6) together with the property (3.13) in Lemma 3.2 to get
$\lim_{N\rightarrow\infty}\frac{1}{N}\log J_{N}^{1}=-\infty,$
which completes the proof of this lemma. ∎
Lemma 3.4.
If $\xi(x)$ is a $pure$-$like$ mixture, then for any $\delta>0$ fixed, there exists some $\varepsilon>0$ sufficiently small such that
$\lim_{N\rightarrow\infty}\mathbb{E}\mathcal{B}_{N,\xi}^{\varepsilon,\delta}=0.$ (3.20)
Proof.
To prove this lemma, it suffices to prove that the equality (3) converges to 00 as $N$ tends to $\infty$ with the event $A_{N}^{\xi}$ replaced by the event
$B_{N}^{\xi}=\left\{\left|\lambda_{\min}\left(\mathrm{GOE}_{N-1}\right)-\frac{\sqrt{N}(y+y_{0})}{\sqrt{(N-1)\xi^{\prime\prime}}}+\frac{2\sqrt{N}}{\sqrt{N-1}}\right|>\frac{\sqrt{N}\delta}{\sqrt{(N-1)p(p-1)}}\right\}.$
Following the same partition with respect to variable $y$ as in (3), we denote the three parts by $I_{N}^{1}$, $I_{N}^{2}$ and $I_{N}^{3}$, respectively. For sufficiently small $\varepsilon$, by using the same argument as in (3.19), we have
$\lim_{N\rightarrow\infty}(I_{N}^{2}+I_{N}^{3})=0.$ (3.21)
For $I_{N}^{1}$, since $y\in(y_{0}-\alpha\sqrt{\varepsilon},y_{0}+\alpha\sqrt{\varepsilon})$, once the event $B_{N}^{\xi}$ occurs, for any $\delta>0$, we take $\varepsilon>0$ to be sufficiently small and $N$ sufficiently large such that
$\left|\lambda_{\min}\left(\mathrm{GOE}_{N-1}\right)+2\right|>\frac{\delta}{2\sqrt{\xi^{\prime\prime}}}.$
Then we deduce for any $\varepsilon$ small enough and $N$ large enough
$I_{N}^{1}$ $\leq C_{\xi,N}\int_{-E_{0}-\varepsilon}^{-E_{0}+\varepsilon}\int_{-y_{0}-\alpha\sqrt{\varepsilon}}^{-y_{0}+\alpha\sqrt{\varepsilon}}\exp\left\{-\frac{N}{2}(x,y)\Sigma^{-1}(x,y)^{\top}\right\}$
×$\quad\times\mathbb{E}\left\{\left|\operatorname{det}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)\xi^{\prime\prime}}}\mathbf{I}_{N-1}\right)\right|1_{\left\{\left|\lambda_{\min}\left(\mathrm{GOE}_{N-1}\right)+2\right|>\frac{\delta}{2\sqrt{\xi^{\prime\prime}}}\right\}}\right\}\mathrm{d}x\mathrm{d}y.$
Combining the Cauchy–Schwarz inequality, second moment upper bound (2) and the LDP for smallest eigenvalues of GOE matrices (2.24), the logarithmic asymptotics formula (3.6) implies that there exists a constant $C>0$ such that for all $\varepsilon>0$ sufficiently small and $N$ large enough
$I_{N}^{1}\leq\exp\{-CN\}\exp\left\{\sup_{\begin{subarray}{c}x\in(-E_{0}-\varepsilon,-E_{0}+\varepsilon)\\ y\in(-y_{0}-\alpha\sqrt{\varepsilon},-y_{0}+\alpha\sqrt{\varepsilon})\end{subarray}}NF(x,y)+o(N)\right\}.$
By the key observation (3.13) in Lemma 3.2, for sufficiently small $\varepsilon$, it follows that $\lim_{N\rightarrow\infty}I_{N}^{1}=0$, which together with (3.21) finishes the lemma. ∎
Finally, we finish the proof of Theorem 3.1 with the help of Lemmas 3.3 and 3.4.
Proof of Theorem 3.1.
Note that $GS=-E_{0}$ as given in (3.8) for the $pure$-$like$ mixture models under Assumption I. The proof of this theorem follows a similar argument as in Theorem 2.1. We only need to replace the parameters $c_{p}$ by $y_{0}$ and $r_{p}$ by $2\sqrt{\xi^{\prime\prime}}$, since we can use Lemmas 3.3 and 3.4 to play the role of Lemmas 2.2 and 2.3 in Theorem 2.1, respectively. ∎
Acknowledgements.
Hao Xu would like to thank his previous supervisor Qiang Zeng for introducing us to consider the problem of Hessian spectrum at the global minimum of spherical mixed $p$-spin models. His careful reading and numerous suggestions have greatly improved the quality of this paper.